Logiweb bibliography of test

Up Help

[0] test

[1] base

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-06-20.UTC:12:52:21.411062 = MJD-53906.TAI:12:52:54.411062 = LGT-4657524774411062e-6